At stp one litre of a gas weighs 1.25 grams.the gas contains 85.71%of carbon and 14.29% of hydrogen .the formula of the compund is??

Grade:11

1 Answers

Pawan
6 years ago
Let no of carbon atoms be XAnd no of hydrogen atoms be yAtomic mass of the compound is 12x+yAt stp 22.4ltr of gas weighs 12x+y g1 ltr weighs 1.24gSo molar mass is 22.4×1.24=27.77gNo of carbon atoms=(85.71×27.77)÷1200=2(approx.)No of hydrogen atoms=(14.29×27.77)÷100=4(approx)Hence the formula for the compound is C2H4